A 34 kg block is sliding on a frictionless surface at a rate of 25 m/s. It enters a surface that has a coefficient of kinetic friction of .35. How long does it take the block to stop and how far does it slide before it comes to a stop after it enters the surface with friction?
1
Expert's answer
2020-11-16T06:35:55-0500
"t=\frac{v}{\mu g}=\frac{25}{0.35\cdot 9.8}=7.3\ s"

"2\mu gs=v^2\\2(0.35)(9.8)s=25^2\\s=91\ m"